OpenStudy (anonymous):

Show all working to proving these Trig Identities. (For the sake of learning) 1.) CosA / SecA - TanA = 1 + SinA 2.) Cosx - Cos3x / sin3x - sinx = tan2x

OpenStudy (anonymous):

I dont think 1) is true

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

written funny

OpenStudy (anonymous):

\[\frac{ \cos x }{ \sec x-\tan x} = 1+\sin x\]

OpenStudy (anonymous):

Oh...... i thought it was